# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

### v1

@@ 5.2.2. VLMs-based BTB
Foundational text classification (de Masson D’Autume et al., 2019) 115,000 - 7,600 122,600 Hard and Blurry News classification, sentiment analysis, Wikipedia article classification, and question-and-answer categorization English 5 tasks
MBPA++ (de Masson D’Autume et al., 2019) 881,000 35,000 38,000 954,000 Hard and Blurry News classification, sentiment analysis, Wikipedia article classification, questions and answers categorization, question answering English 9 tasks
Lifelong FewRel (Holla et al., 2020) - - - - Hard and Blurry few-shot relation detection English 10 tasks
Firehose (Hu et al., 2022) - - - 110,000,000 Blurry Personalized online language learning English 1 tasks
TemporalWiki (Jang et al., 2022a) - - - - - - English -
@@ 6.1. Offline Datasets for NLP
The most typical task for continual learning is text classification. The foundational text classification benchmark encompasses five text classification datasets introduced by (Zhang et al., 2015), including AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers (Sun et al., 2020a). Particularly, the AG News dataset has 4 classes for news classification; the Amazon and Yelp dataset has 5 classes for sentiment analysis; the DBpedia dataset has 14 classes for Wikipedia text classification; and the Yahoo dataset has 10 classes for Q&A classification. The text classification benchmark includes 115,000 training and 7,600 test examples for each task, holding out 500 samples per class from the training set for validation. Building upon this, Razdaibiedina et al. (Razdaibiedina et al., 2023) developed a novel continual learning (CL) benchmark. This benchmark not only utilizes the foundational text classification benchmark but also integrates additional datasets from the GLUE benchmark (Wang et al., 2018), SuperGLUE benchmark (Wang et al., 2019a), and the IMDB dataset (Maas et al., 2011). Specifically, the GLUE benchmark datasets included are MNLI, QQP, RTE, and SST2, focusing on tasks like natural language inference, paraphrase detection, and sentiment analysis. Similarly, the SuperGLUE datasets—WiC, CB, COPA, MultiRC, and BoolQ—encompass tasks ranging from word sense disambiguation to question answering. DE&E (Wójcik et al., 2023) uses three common text classification data sets with different characteristics-News-groups, BBC News, and Consumer Finance Complaints2. Such datasets can be used to evaluate the models on tasks with different difficulty levels.
The datasets introduced in (Wang et al., 2023c) further are categorized into two groups based on the domain relevance between tasks: far-domain and near-domain. The far-domain group comprises two text classification tasks, which are foundational benchmarks (Zhang et al., 2015) divided into topic classification (AG News, Yahoo Answers, DBpedia) and sentiment classification (Yelp, Amazon Reviews). In contrast, the near-domain group uses the Web of Science (WOS) (Kowsari et al., 2017) and 20 Newsgroups (Lang, 1995), which are restructured according to their high inter-task relevance. The WOS dataset comprises seven parent classes, each with five closely related sub-classes, while the 20 Newsgroups dataset, containing six news topics, is reorganized into four tasks to maximize inter-task correlation.
@@
Some studies focus on intent classification tasks, where the classes are quite different in various domains or scenarios. In the realm of intent classification and detection, several datasets have been specifically designed to advance the field by addressing different challenges and providing diverse environments for model training and evaluation. The dataset, as introduced in PAGeR (Varshney et al., 2022), aims to tackle the lifelong intent detection problem by combining three public intent classification datasets (CLINC150 (Larson et al., 2019), HWU64 (Liu et al., 2021a), BANKING77 (Casanueva et al., 2020)), one text classification dataset (Stackoverflow S20 (Xu et al., 2017)), and two public multidomain dialog intent detection datasets (SGD (Rastogi et al., 2020), MWOZ (Budzianowski et al., 2018)). Moreover, FewRel (Han et al., 2018) is also incorporated to tackle the lifelong relation extraction problem. This integration is intended to simulate real-world applications by encompassing a broad spectrum of domains and query distributions, thereby facilitating the development of more robust and versatile intent detection systems.
Conversely, the dataset compiled in PLE (Li et al., 2022b) consolidates nine well-regarded intent detection datasets, including CLINC150 (Larson et al., 2019) and HWU64 (Liu et al., 2021a), among others, arranged in a fixed random sequence to form a standardized benchmark. This dataset emphasizes the importance of consistency and comparability in performance evaluations across different intent detection models, providing a platform for assessing and enhancing various methodologies. The dataset described by MeLL (Wang et al., 2021a) specifically addresses intent detection within two distinct contexts: task-oriented dialogues (TaskDialog-EUIC) and real-world e-commerce interactions (Hotline-EUIC). TaskDialog-EUIC integrates data from Snips (Coucke et al., 2018), TOP semantic parsing (Gupta et al., 2018), and Facebook’s Multilingual Task Oriented Dataset (Schuster et al., 2018) into 90 tasks with overlapping label sets, amounting to over ten thousand samples. Hotline-EUIC is derived from an e-commerce dialogue system (Li et al., 2017) and the hotline audios are transcribed to text by a high-accuracy industrial Automatic Speech Recognition (ASR) system.
@@
Ke et al. (Ke and Xu, 2021) developed a task incremental learning dataset for aspect-based sentiment classification (ABSC). This dataset aggregates reviews from four distinct sources, thereby enhancing its diversity and applicability across multiple domains. The sources include the L5Domains dataset by Hu et al. (Hu and Liu, 2004), which features consumer reviews for five different products; the Liu3Domains dataset by Liu (Liu et al., 2015), comprising reviews pertaining to three products; the Ding9Domains dataset by Ding et al. (Ding et al., 2008), which includes reviews of nine varied products; and the SemEval14 dataset, which is focused on reviews of two specific products—laptops and restaurants.
@@
In the rapidly advancing field of machine learning, diverse datasets function as crucial benchmarks for exploring various dimensions of language and code generation. These datasets address both universal and task-specific challenges, enabling a comprehensive evaluation of model capabilities. A particularly significant dataset highlighted in the work by Continual-T0 (Scialom et al., 2022b) focuses on English language generation tasks, including text simplification and empathetic dialogue generation, among others (Camburu et al., 2018; Bin Tareaf, 2017). The design of this dataset maintains uniformity in size, facilitating effective comparative analyses of performance across distinct tasks by ensuring a consistent volume of data for training. In a subsequent study, Luo et al. (Luo et al., 2023) conduct an analysis of catastrophic forgetting on Bloomz (Scao et al., 2022) using Continual T0 datasets.
@@
The dataset, introduced in LAMOL (Sun et al., 2020a), integrates elements from both DecaNLP (McCann et al., [n. d.]) and the foundational text classification benchmark (Zhang et al., 2015). This dataset encompasses five distinct NLP tasks originally sourced from DecaNLP: question answering, semantic parsing, sentiment analysis, semantic role labeling, and goal-oriented dialogue. For the purposes of this dataset, all tasks, whether derived from DecaNLP or the foundational text classification benchmark, are restructured into a uniform format, conceptualized under the framework of a question answering task. Moreover, the dataset devised in RVAE_LAMOL (Wang et al., 2022a), employs three tasks from DecaNLP: the English Wizard of Oz (WOZ) for goal-oriented dialogue, QA-SRL for semantic role labeling in a SQuAD-style format, and SST, which is a binary version of the Stanford Sentiment Treebank categorizing sentiments as positive or negative. These tasks are specifically treated as sequence generation tasks.
@@
The dataset introduced in COPR (Zhang et al., 2023b) represents a pioneering effort in applying both Task Incremental Learning (TIL) and Domain Incremental Learning (DIL) within the context of benchmarks that utilize existing human preferences. Specifically, the TIL framework in this dataset mandates that the model sequentially acquires knowledge from three distinct tasks. These include the question-answering task utilizing the HH-RLHF dataset (Bai et al., 2022), the summarization task based on the Reddit TL, DR dataset with human feedback (Volske et al., 2017), and the positive film review generation task using the IMDB dataset (Maas et al., 2011). Meanwhile, the DIL framework requires the model to adapt to three distinct segments from the SHP dataset, as described by Ethayarajh et al. (Ethayarajh et al., 2022).
@@
The dataset described in Adaptive Compositional Modules (Zhang et al., 2022b) explores sequence generation and categorizes tasks into ”similar” and ”dissimilar” groups based on their characteristics. Tasks classified as similar, including E2ENLG (Novikova et al., 2017) and four domains (restaurant, hotel, TV, laptop) from RNNLG (Wen et al., 2015), demonstrate shared patterns and are tested across four sequence orders, comprising a total of five tasks. In contrast, dissimilar tasks such as WikiSQL (SQL query generation) (Zhong et al., 2017), CNN/DailyMail (news article summarization) (See et al., 2017), and MultiWOZ (semantic state sequence generation) (Budzianowski et al., 2018) exhibit significant distributional shifts from previously encountered tasks. The CODETASKCL dataset, explored by Yadav et al. (Yadav et al., 2023), encompasses a diverse array of code-centric tasks, including code generation (Iyer et al., 2018), summarization (Husain et al., 2019), translation (Lu et al., 2021), and refinement (Tufano et al., 2019) across various programming languages. This dataset significantly enhances the breadth of language processing applications within technical fields.
@@
In the realm of natural language processing (NLP), various datasets are tailored to specific aspects of the task under continual learning paradigms. The dataset introduced in ExtendNER (Monaikul et al., 2021), exemplifies a continual learning approach to Named Entity Recognition (NER). This dataset amalgamates the CoNLL-03 English NER (Sang and De Meulder, 2003) and OntoNotes (Hovy et al., 2006), covering a broad spectrum of entity types and sources. This hybrid dataset is structured to challenge the adaptability and generalization capabilities of NER systems across varied contexts. Unlike the static nature of text in NER tasks, the Schema-Guided Dialog (SGD) (Rastogi et al., 2020) dataset, utilized in C-PT (Zhu et al., 2022), serves the Dialog State Tracking aspect of IE, which involves maintaining the context of a dialog over time. The SGD dataset features 44 services across 19 domains, each treated as a separate task, and is designed to evaluate models on their ability to manage and extract information across conversational turns. Lastly, the lifelong SimpleQuestions and lifelong FewRel datasets, devised in (Wang et al., 2019b) is crafted for the task of relation extraction. It merges elements from the SimpleQuestions (Bordes et al., 2015) and FewRel (Han et al., 2018) to form a lifelong learning benchmark that confronts the challenges of relation detection in a few-shot context.
@@
In the realm of continual pre-training for large language models (LMs), the development and utilization of specialized benchmarks play a pivotal role in evaluating and enhancing the effectiveness of continual learning systems. The dataset, introduced in CPT (Ke et al., 2022), primarily focuses on the continual post-training of LMs across a series of domain-specific, unlabeled datasets. It provides a rigorous test environment by using diverse corpora such as Yelp Restaurant Reviews (Xu et al., 2019), AI and ACL Papers (Lo et al., 2020), and AGNews articles (Zhang et al., 2015). Its main objective is to gauge how well an LM can incrementally integrate domain-specific knowledge without forgetting previously learned information, thereby enhancing its few-shot learning capabilities in these domains. Contrary to the datasets employed in CPT (Ke et al., 2022), which evaluate domain-specific adaptability and incremental learning, the CKL benchmark (Jang et al., 2022b) is meticulously designed to measure the LM’s ability to retain timeless knowledge, update obsolete information, and acquire new knowledge. It comprises subsets like INVARIANTLAMA, UPDATEDLAMA, and NEWLAMA, which are crafted to probe specific types of knowledge that an LM may encounter in its learning trajectory.
Whereas the aforementioned two datasets assess more controlled dimensions of knowledge integration and retention, the dataset introduced in ELLE (Qin et al., 2022) focuses on the dynamic scenario of accumulating streaming data from diverse sources in a lifelong learning context. This dataset mirrors the real-world challenge of a language model (LM) that must continuously adapt to new data inflows from multiple domains, including BOOKCORPUS (WB) (Zhu et al., 2015), NEWS ARTICLES (NS) (Zellers et al., 2019), AMAZON REVIEWS (REV) (He and McAuley, 2016), BIOMEDICAL PAPERS (BIO) (Lo et al., 2020) and COMPUTER SCIENCE PAPERS (CS) (Lo et al., 2020). The benchmark evaluates the LM’s capacity to effectively integrate new information from these varied sources over time, highlighting the essential need for LMs to evolve in response to continual data growth and shifts in data distribution. Jin et al. (Jin et al., 2022) construct data streams to represent two prevalent types of domain shifts observed in practical scenarios. The first, a Domain-incremental Paper Stream, simulates the sequential evolution of research areas within academic papers, encompassing diverse disciplines such as biomedical and computer science. The second, a Chronologically-ordered Tweet Stream, models the temporal progression of tweets over time.
@@
An increasing number of datasets are adopting a hybrid task approach that integrates multiple learning paradigms and task types, aimed at testing and enhancing the adaptability of models. A notable example is the dataset introduced in AdapterCL (Madotto et al., 2020), which is tailored for task-oriented dialogue systems. This dataset incorporates four task-oriented datasets: TaskMaster 2019 (TM19) (Byrne et al., 2019), TaskMaster 2020 (TM20) (Byrne et al., 2019), Schema Guided Dialogue (SGD) (Rastogi et al., 2020), and MultiWoZ (Budzianowski et al., 2018). These datasets have been pre-processed to form a curriculum encompassing 37 domains, structured under four continual learning settings: INTENT classification, Dialogue State Tracking (DST), Natural Language Generation (NLG), and end-to-end (E2E) modeling.
Continual Instruction Tuning Benchmark (CITB) (Zhang et al., 2023a) extends the concept of continual learning by focusing on instruction-based NLP tasks. Built on the comprehensive SuperNI (Wang et al., 2022c) dataset, it includes over 1,600 tasks across diverse NLP categories. CITB differentiates itself by formulating two distinct streams—InstrDialog and InstrDialog++—to examine how models integrate and retain new dialogue-oriented and varied NLP tasks under continual learning settings. This benchmark suite not only tests task retention and adaptability but also explores how instruction tuning can be optimized for a continual learning framework. The ConTinTin (Yin et al., 2022) is an adaptation of the NATURAL-INSTRUCTIONS dataset, specifically restructured to facilitate a continual learning framework. This adaptation involves decomposing the original crowdsourcing instructions into smaller, distinct sub-tasks to create a new dataset. Additionally, the new dataset incorporates a novel experimental design where tasks are selected randomly to create diverse sequences, enabling the evaluation of a model’s adaptability to novel instructions without prior exposure.
The dataset, used in Conure (Yuan et al., 2021), consists of Tencent TL (TTL) (Yuan et al., 2020) and Movielens (ML). The TTL dataset is designed to address three item recommendation tasks and three user profiling tasks, whereas the ML dataset exclusively focuses on three item recommendation tasks. Both datasets have been pre-processed to facilitate a continual learning framework, simulating environments where models must adapt to evolving data streams. Furthermore, Kim et al. (Kim et al., 2023) introduced the proprietary NAVER Shopping dataset, which builds upon the previously mentioned datasets. The NAVER Shopping dataset features six tasks: two for search query prediction, two for purchased item category prediction, and two for user profiling, all designed to meet real-world industry requirements.
Finally, the TRACE dataset, introduced by Wang et al. (Wang et al., 2024a), is specifically designed to bridge the existing gap in the evaluation of large language models (LLMs) within the continual learning framework, encompassing a wide range of complex and specialized tasks. Distinguished from other datasets, TRACE targets domain-specific tasks that are multilingual and technical, including code completion and mathematical reasoning. This diversity presents a unique set of challenges that span both specialized and broad dimensions. Moreover, TRACE rigorously assesses the models’ ability to sustain performance across tasks that demand different knowledge bases and cognitive skills. This evaluation highlights the essential need for adaptability in LLMs trained to operate under continual learning conditions, underscoring their potential in dynamic real-world applications.

### v2

@@ 6.2.2. VLMs-based BTB
MBPA++ (de Masson D’Autume et al., 2019) 115,000 - 7,600 122,600 Hard and Blurry Mixed 5 classification tasks AGNews, Amazon Reviews, Yelp Reviews, DBpedia, Yahoo Answers English 5 tasks
MBPA++ (de Masson D’Autume et al., 2019) 306,000 35,000 - 341,000 Hard and Blurry Question answering QuAC, SQuAD, Trivia Wikipedia, Trivia Web English 4 tasks
Lifelong FewRel (Holla et al., 2020) - - - - Hard and Blurry Few-shot relation detection FewRel English 10 tasks
Firehose (Hu et al., 2022) - - - 110,000,000 Blurry Personalized online language learning Twitter English 1 tasks
TemporalWiki (Jang et al., 2022a) - - - - - - Wikipedia English -
@@ 7. Datasets
In this section, we review the typical datasets for offline and online continual learning for different tasks (Table 1).
@@ 7.1. Offline Datasets for NLP
The most typical task for continual learning is text classification. The foundational text classification benchmark encompasses five text classification datasets introduced by (Zhang et al., 2015), including AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers (Sun et al., 2020a). Particularly, the AG News dataset has 4 classes for news classification; the Amazon and Yelp dataset has 5 classes for sentiment analysis; the DBpedia dataset has 14 classes for Wikipedia text classification; and the Yahoo dataset has 10 classes for Q&A classification. Building upon this, Razdaibiedina et al. (Razdaibiedina et al., 2023) developed a novel continual learning (CL) benchmark. This benchmark not only utilizes the foundational text classification benchmark but also integrates additional datasets from the GLUE benchmark (Wang et al., 2018), SuperGLUE benchmark (Wang et al., 2019a), and the IMDB dataset (Maas et al., 2011). DE&E (Wójcik et al., 2023) uses three common text classification data sets with different characteristics-News-groups, BBC News, and Consumer Finance Complaints2. Such datasets can be used to evaluate the models on tasks with different difficulty levels.
The datasets introduced in (Wang et al., 2023c) further are categorized into two groups based on the domain relevance between tasks: far-domain and near-domain. The far-domain group comprises two text classification tasks, which are foundational benchmarks (Zhang et al., 2015) divided into topic classification (AG News, Yahoo Answers, DBpedia) and sentiment classification (Yelp, Amazon Reviews). In contrast, the near-domain group uses the Web of Science (WOS) (Kowsari et al., 2017) and 20 Newsgroups (Lang, 1995), which are restructured according to their high inter-task relevance.
@@
Some studies focus on intent classification tasks, where the classes are quite different in various domains or scenarios. The dataset, as introduced in PAGeR (Varshney et al., 2022), aims to tackle the lifelong intent detection problem by combining three public intent classification datasets (CLINC150 (Larson et al., 2019), HWU64 (Liu et al., 2021a), BANKING77 (Casanueva et al., 2020)), one text classification dataset (Stackoverflow S20 (Xu et al., 2017)), and two public multidomain dialog intent detection datasets (SGD (Rastogi et al., 2020), MWOZ (Budzianowski et al., 2018)). Moreover, FewRel (Han et al., 2018) is also incorporated to tackle the lifelong relation extraction problem. This integration is intended to simulate real-world applications by encompassing a broad spectrum of domains and query distributions, thereby facilitating the development of more robust and versatile intent detection systems.
Conversely, the dataset compiled in PLE (Li et al., 2022b) consolidates nine well-regarded intent detection datasets, including CLINC150 (Larson et al., 2019) and HWU64 (Liu et al., 2021a), among others, arranged in a fixed random sequence to form a standardized benchmark. The dataset described by MeLL (Wang et al., 2021) specifically addresses intent detection within two distinct contexts: task-oriented dialogues (TaskDialog-EUIC) and real-world e-commerce interactions (Hotline-EUIC). TaskDialog-EUIC integrates data from Snips (Coucke et al., 2018), TOP semantic parsing (Gupta et al., 2018), and Facebook’s Multilingual Task Oriented Dataset (Schuster et al., 2018) into 90 tasks with overlapping label sets, amounting to over ten thousand samples.
@@
Ke et al. (Ke and Xu, 2021) develop a task incremental learning dataset for aspect-based sentiment classification (ABSC). This dataset aggregates reviews from four distinct sources, thereby enhancing its diversity and applicability across multiple domains. The sources include the L5Domains dataset by Hu et al. (Hu and Liu, 2004), the Liu3Domains dataset (Liu et al., 2015), the Ding9Domains dataset (Ding et al., 2008), and the SemEval14 dataset.
@@
Diverse datasets function as crucial benchmarks for exploring various dimensions of language and code generation. A particularly significant dataset highlighted in the work by Continual-T0 (Scialom et al., 2022a) focuses on English language generation tasks, including text simplification and empathetic dialogue generation, among others (Camburu et al., 2018; Bin Tareaf, 2017). In a subsequent study, Luo et al. (Luo et al., 2023) conduct an analysis of catastrophic forgetting on Bloomz (Scao et al., 2022) using Continual T0 datasets. The dataset, introduced in LAMOL (Sun et al., 2020a), integrates elements from both DecaNLP (McCann et al., [n. d.]) and the foundational text classification benchmark (Zhang et al., 2015). Moreover, the dataset devised in RVAE_LAMOL (Wang et al., 2022a), employs three tasks from DecaNLP: the English Wizard of Oz (WOZ) for goal-oriented dialogue, QA-SRL for semantic role labeling in a SQuAD-style format, and SST, which is a binary version of the Stanford Sentiment Treebank.
@@
The dataset introduced in COPF (Zhang et al., 2023b) represents a pioneering effort in applying both TIL and DIL within the context of benchmarks that utilize existing human preferences. Specifically, the TIL framework in this dataset mandates that the model sequentially acquires knowledge from three distinct tasks. These include the question-answering task utilizing the HH-RLHF dataset (Bai et al., 2022), the summarization task based on the Reddit TL, DR dataset with human feedback (Volske et al., 2017), and the positive film review generation task using the IMDB dataset (Maas et al., 2011). Meanwhile, the DIL framework requires the model to adapt to three distinct segments from the SHP dataset, as described by Ethayarajh et al. (Ethayarajh et al., 2022).
@@
The dataset described in ACM (Zhang et al., 2022) explores sequence generation and categorizes tasks into “similar” and “dissimilar” groups based on their characteristics. Tasks classified as similar, including E2ENLG (Novikova et al., 2017) and four domains (restaurant, hotel, TV, laptop) from RNNLG (Wen et al., 2015), demonstrate shared patterns and are tested across four sequence orders, comprising a total of five tasks. In contrast, dissimilar tasks such as WikiSQL (Zhong et al., 2017), CNN/DailyMail (See et al., 2017), and MultiWOZ (Budzianowski et al., 2018) exhibit significant distributional shifts from previously encountered tasks. The CODETASKCL dataset, explored by Yadav et al. (Yadav et al., 2023), encompasses a diverse array of code-centric tasks, including code generation (Iyer et al., 2018), summarization (Husain et al., 2019), translation (Lu et al., 2021), and refinement (Tufano et al., 2019) across various programming languages. This dataset significantly enhances the breadth of language processing applications within technical fields.
@@
The dataset introduced in ExtendNER (Monaikul et al., 2021), exemplifies a continual learning approach to Named Entity Recognition (NER). This dataset amalgamates the CoNLL-03 English NER (Sang and De Meulder, 2003) and OntoNotes (Hovy et al., 2006), covering a broad spectrum of entity types and sources. This hybrid dataset is structured to challenge the adaptability and generalization capabilities of NER systems across varied contexts. Unlike the static nature of text in NER tasks, the Schema-Guided Dialog (SGD) (Rastogi et al., 2020) dataset, utilized in C-PT (Zhu et al., 2022), serves the Dialog State Tracking aspect of IE, which involves maintaining the context of a dialog over time. The SGD dataset features 44 services across 19 domains, each treated as a separate task, and is designed to evaluate models on their ability to manage and extract information across conversational turns. Lastly, the lifelong SimpleQuestions and lifelong FewRel datasets, devised in (Wang et al., 2019b) is crafted for the task of relation extraction. It merges elements from the SimpleQuestions (Bordes et al., 2015) and FewRel (Han et al., 2018) to form a lifelong learning benchmark that confronts the challenges of relation detection in a few-shot context.
@@
In the realm of continual pre-training for LLMs, the development and utilization of specialized benchmarks play a pivotal role in evaluating and enhancing the effectiveness of continual learning systems. The dataset, introduced in CPT (Ke et al., 2022), primarily focuses on the continual post-training of LMs across a series of domain-specific, unlabeled datasets. It provides a rigorous test environment by using diverse corpora such as Yelp Restaurant Reviews (Xu et al., 2019), AI and ACL Papers (Lo et al., 2020), and AGNews articles (Zhang et al., 2015). Its main objective is to gauge how well an LM can incrementally integrate domain-specific knowledge without forgetting previously learned information, thereby enhancing its few-shot learning capabilities in these domains. Contrary to the datasets employed in CPT (Ke et al., 2022), which evaluate domain-specific adaptability and incremental learning, the CKL benchmark (Jang et al., 2022b) is meticulously designed to measure the LM’s ability to retain timeless knowledge, update obsolete information, and acquire new knowledge. It comprises subsets like INVARIANTLAMA, UPDATEDLAMA, and NEWLAMA, which are crafted to probe specific types of knowledge that an LM may encounter in its learning trajectory.
Whereas the aforementioned two datasets assess more controlled dimensions of knowledge integration and retention, the dataset introduced in ELLE (Qin et al., 2022) focuses on the dynamic scenario of accumulating streaming data from diverse sources in a lifelong learning context. This dataset mirrors the real-world challenge of a LM that must continuously adapt to new data inflows from multiple domains, including BOOKCORPUS (Zhu et al., 2015), NEWS ARTICLES (Zellers et al., 2019), AMAZON REVIEWS (He and McAuley, 2016), BIOMEDICAL PAPERS (Lo et al., 2020) and COMPUTER SCIENCE PAPERS (Lo et al., 2020). The benchmark evaluates the LM’s capacity to effectively integrate new information from these varied sources over time, highlighting the essential need for LMs to evolve in response to continual data growth and shifts in data distribution. Jin et al. (Jin et al., 2022) construct data streams to represent two prevalent types of domain shifts observed in practical scenarios. The first, a Domain-incremental Paper Stream, simulates the sequential evolution of research areas within academic papers, encompassing diverse disciplines such as biomedical and computer science. The second, a Chronologically-ordered Tweet Stream, models the temporal progression of tweets over time.
@@
An increasing number of datasets are adopting a hybrid task approach that integrates multiple learning paradigms and task types, aimed at enhancing the adaptability of models. AdapterCL (Madotto et al., 2020) introduces a dataset for task-oriented dialogue systems, which incorporates four task-oriented datasets: TaskMaster 2019 (Byrne et al., 2019), TaskMaster 2020 (Byrne et al., 2019), Schema Guided Dialogue (Rastogi et al., 2020), and MultiWoZ (Budzianowski et al., 2018). Continual Instruction Tuning Benchmark (CITB) (Zhang et al., 2023a) extends the concept of continual learning by focusing on instruction-based NLP tasks. Built on the comprehensive SuperNI (Wang et al., 2022c) dataset, it includes over 1,600 tasks across diverse NLP categories. The ConTinTin (Yin et al., 2022) is an adaptation of the NATURAL-INSTRUCTIONS dataset, specifically restructured to facilitate a CL framework.
The dataset, used in Conure (Yuan et al., 2021), consists of Tencent TL (TTL) (Yuan et al., 2020) and Movielens (ML). The TTL dataset is designed to address three item recommendation tasks and three user profiling tasks, whereas the ML dataset exclusively focuses on three item recommendation tasks. Furthermore, Kim et al. (Kim et al., 2023) introduced the proprietary NAVER Shopping dataset, which builds upon the previously mentioned datasets. The NAVER Shopping dataset features six tasks: two for search query prediction, two for purchased item category prediction, and two for user profiling, all designed to meet real-world industry requirements. Finally, the TRACE dataset, introduced by Wang et al. (Wang et al., 2024b), is specifically designed to bridge the existing gap in the evaluation of LLMs within the CL framework, encompassing a wide range of complex and specialized tasks.